what is larmor radius for electron and ion?
1
Expert's answer
2020-08-06T16:46:20-0400

The Larmor radius for an electron with a mass "m_e" and charge "e" moving in a magnetic field "B" perpendicular to the velocity of electron "v" can be calculate as


"r=\frac{m_ev}{eB},"


substituting the numerical values, we get


"r(v,B)=\frac{5.686\cdot10^{-12}\cdot v}{B}."

For an ion of mass "m" and charge "q" , the same equation can be applied under the conditions stated above:


"r=\frac{mv}{qB}."